40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11395     Entry time: Wed Jul 8 17:46:20 2015
Author: Jessica 
Type: Summary 
Category: General 
Subject: Updated Time Delay Plots 

I re-measured the transfer function for Cable B, because the residuals in my previous post for cable B indicated a bad fit. 

I also realized I had made a mistake in calculating the time delay, and calculated more reasonable time delays today. 

Cable A had a delay of 202.43 +- 0.01 ns.

Cable B had a delay of 202.44 +- 0.01 ns. 

Attachment 1: resid_CableA.png  75 kB  | Hide | Hide all
resid_CableA.png
Attachment 2: resid_CableB.png  83 kB  | Hide | Hide all
resid_CableB.png
ELOG V3.1.3-